OpenStudy (anonymous):

Show that \(\log_23\) is irrational. Proof: Let \(\log_23\) be a rational number. Then \(\log_23=a/b\), where \((a,b)=1\) with \(b\neq0\). It follows that \(2^{a/b}=3\)... I get stuck there. How could I proceed?

OpenStudy (anonymous):

Perhaps: ... \(2^a=3^b\)?

OpenStudy (anonymous):

I see how I can reach a contradiction now.

OpenStudy (anonymous):

Yeah, and \(2^a = 3^b, b \neq 0\) isn't possible

OpenStudy (anonymous):

So, by contradiction \(\log_23\) is rational

OpenStudy (anonymous):

irrational*